OpenStudy (anonymous):

find the linearization of f(x)=cosx at x=pi/2

OpenStudy (anonymous):

Find f ' (pi/2), that is the slope of the linearization. Then you need the point (pi/2, f(pi/2)) Then write an equation in point-slope form f'(x) = -sin x What is the -sin (pi/2)? Find cos(pi/2) Is this enough of a how to?

OpenStudy (anonymous):

this = f(pi/2)+f'(pi/2)(x-pi/2) if i remember the formula correctly = cos (pi/2) + -sine pi/2(x - pi/2) = 0 - 1(x - pi/2)

OpenStudy (anonymous):

finalizing jimyrep's answer \[\pi/2 - x\]